Circular No. 6998

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998dq IN NGC 6754 Brett White, Linden, New South Wales, has reported — directly and

via R. Evans (Hazelbrook, N.S.W.) and B. Schmidt (Mount Stromlo and Siding Spring Observatories) — his visual discovery with a 0.41-m reflector on Aug. 23.4 UT of an apparent supernova located ∼ 18′′ west of the nucleus of NGC 6754; the visual magnitude of the new star was reported by Evans as being 14.3. An image obtained by T. Dobosz (Nepean Astronomy Centre) on Aug. 24 yielded an offset of 20′′ west. Schmidt reports that BVRI CCD images obtained on the MSSSO 0.76-m telescope by B. Crook on Aug. 23.6 show a new blue object at α = 19^{h}11^{m}23^{s}.78, δ = −50^{o}38′25′′.5 (equinox 2000.0), which is 18′′.7 west and 5′′ north of the galaxy nucleus. SN 1998dq is located ∼ 5′′ from SN 1998X, is clearly a new object, and is extremely blue (B −V ∼ 0). SN 1998dq is also seen on CCD images taken on Aug. 24.6 by A. Gutierrez with the Mount Stromlo 0.76-m telescope. Comparison with the discovery images of SN 1998X (cf. IAUC 6847) shows that SN 1998dq is not present on Mar. 13.9 to a limiting magnitude of R = 18.

XTE J1710–281 = 1RXS J171012.3–280754 C. B. Markwardt, National Research Council and Goddard Space

Flight Center (GSFC); F. E. Marshall and J. Swank, GSFC; and T. Takeshima, Universities Space Research Association and GSFC, report: “A transient x-ray source was serendipitously detected with the RXTE PCA in two slews, first on July 22.0 UT at a level of 24 counts/s, and on Aug. 12.7 at a level of 64 counts/s (per five PCUs). In subsequent PCA scans on Aug. 19.3, the source had decreased to 23 counts/s (5 × 10^{−11} erg s^{−1} cm^{−2} over 2–10 keV). The best-fit position is α = 17^{h}10^{m}12^{s}, δ = −28^{o}07′.1 (equinox 2000.0), with a position uncertainty of ± 2′ in α and ± 4′ in δ (90-percent confidence), and it is consistent with the location of the unidentified source 1RXS J171012.3–280754 of the ROSAT All-Sky Survey Bright Source Catalogue. The spectrum on Aug. 19.3 is consistent with either thermal bremsstrahlung (kT = 14 ± 3 keV) or a power law (photon index 1.8 ± 0.1), with interstellar absorption < 2 × 10^{22} cm^{−2}. No spectral lines are apparent. No short-timescale variability was observed, although the 400-s pointed observation only provided a 3σ upper limit of 14 percent for the best case of coherent oscillations. Further observations at other wavelengths are desirable.”

1998 August 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green